OpenStudy (anonymous):

Can someone please help me? Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f o g)(-5) A. -53 B. 3 C. -9 D. 23

OpenStudy (anonymous):

f o g(x) = f(g(x))

OpenStudy (anonymous):

that is f(-4x+3)

OpenStudy (anonymous):

that is -2(-4x+3) - 7 that is = 8x-6-7 that is = 8x-13

OpenStudy (anonymous):

at x= -5, it becomes 8(-5)-13 that = -40-13 = -53

OpenStudy (anonymous):

answer is A

OpenStudy (anonymous):

Thank You!